# Epidemic Forecasting with a Hybrid Deep Learning Method Using CNN-LSTM With WOA-GWO Parameter Optimization: Global COVID-19 Case Study

### v2

@@ 1 Introduction
COVID-19 affected the lives of people all over the world and took the lives of thousands around the world. On the other hand, the social and economic consequences of COVID-19 sometimes inflicted irreparable damage on countries, and it was able to confront the world with new challenges and crises. Therefore, many researchers have tried to minimize the damage caused by this disease by providing models to predict the rate of coronary artery disease. Various approaches to forecasting have also been developed. For instance, [He2020ADS] developed a discrete-time stochastic model to describe the dynamics of pandemic expansion. Prediction Models Traditional time series have been routinely studied to predict COVID-19 cases [kucharski2020early, wu2020nowcasting, zhuang2020preliminary]. However, Stochastic modeling has some drawbacks such as being incapable of handling large data scales, modeling complex nonlinear characteristics, and its dependence on pre-defined features or assumptions[ihme2020forecasting]. Furthermore, the traditional Autoregressive Integrated Moving Average (ARIMA) model proved to be a suitable approach for describing short-term autocorrelation in time series data, researchers have commonly used it. [collou2019digitalization] utilized a time series method such as the ARIMA to predict the number of validated items and compared to other conventional methods such as stochastic modeling, which ARIMA proved to be more accurate in short-time forecasting. Other studies have been developed in India using traditional ARIMA modeling [gupta2020trend]. In 2022, other researchers from Bangladesh used ARIMA to predict verified people, deaths, and recoveries in COVID-19 [nesa2022forecasting]. [ribeiro2020short] used ARIMA, cubist regression (CUBIST), random forest (RF), Ridge Regression (RIDGE), Support Vector Regression (SVR), and stacking-ensemble learning to forecast the total number of patients in Brazil in 1, 3, and 6 days. ARIMA has had superior performance in the case of short-time forecasting. There are also some other studies devoted to developing ARIMA approaches for forecasting the Covid epidemy [roy2021spatial, kufel2020arima, singh2020development, moftakhar2020exponentially, liu2020predicting, duan2020arima, arunkumar2021forecasting].
@@
While ARIMA may provide a good fit for the initial spread of COVID-19, its effectiveness may diminish as the pandemic progresses and new interventions are implemented [hernandez2020forecasting]. On the other hand, another approach, namely logistic function, may allow for a more flexible and dynamic approach to modeling the pandemic’s impact on a particular population. There are studies on developing a logistic function to predict the spread of COVID-19 [roosa2020real]. The logistic function method assumes that the spread of COVID-19 can be modeled using a sigmoidal curve, which allows for the prediction of an eventual plateau or decline in the number of cases, which are discussed completely in [chen2020reconstructing, li2020trend, qeadan2020naive] which proved to be an appropriate approach when the number of cases is decreasing.
@@
In [wang2020prediction] a more advanced algorithm named Logistic Patient Information Based Algorithm (LPIBA) to estimate the number of COVID-19-related deaths in China. Their results show that the overall mortality rate in Hubei and Wuhan is 13 percent and between 0.75 percent and 3 percent in the rest of China. Furthermore, in [fathollahi2022sustainable], the LPIBA was developed to address forecasting the infected cases. The LPIBA proved to have superior performance over ARIMA, having features such as Incorporation of patient information, flexibility in terms of data type, and ability to handle binary outcomes. However, it is worth mentioning the LPIBA has limitations when detailed data is not available or it is not feasible to collect some features. [mohan2022predicting] used a hybrid ARIMA and Prophet Model to predict daily confirmed and cumulative confirmed cases in India. The Prophet model is open-source developed by Facebook Data and proved to have appropriate performance with time-series types of data having seasonal effects [taylor2018forecasting, taylor2017prophet]. It proved to be a robust algorithm in the case when there are missing data [ndiaye2020analysis]. In another work, Hosseini et al. proposed a Bayesian predictive method to predict the West Nile virus cases[hosseini2024parsimonious], however, these methods also need careful parameter selection and prior distribution assumptions.
@@
Another important and popular approach to epidemic modeling involves treating disease spread as stochastic processes over networks [pastor2015epidemic]. Historically, these methods were constrained to small populations, typically on the order of a few thousand individuals. However, the recently introduced FastGEMF module has enabled simulations encompassing millions of nodes over multi or single-layer complex networks [10876117]. While network-based disease spread models provide precise insights into disease propagation, their stochastic nature and the high degree of freedom (DOF) in selecting network parameters or disease models—such as SIS, SIR, and others—often necessitate extensive simulations. This approach is particularly suitable when substantial information is available regarding disease dynamics or population structure, as demonstrated in studies such as [das2024sir] that model spill-over using the SIR model over interconnected networks with a case study of the Ebola virus to see how the level of interconnection can lead to an epidemic in human networks. In contrast, for newly emerging diseases like COVID-19, purely data-driven methods are often more appropriate [XU2022105342].
@@
To deal with more complex data, Neural Networks proved to have superior performance over Prophet Model and ARIMA[bengio2013representation, schmidhuber2015deep, tygert2016mathematical]. In [de2022covid], authors used the ANN algorithm to forecast the number of cumulative cases of infection and deaths in Brazil. They also used a substantial mitigation procedure adopted (mandatory use of masks) was experimented as an input to evaluate the improvement in the results. [vadyala2021prediction] developed a more complex NN, a hybrid algorithm for predicting short-term COVID-19 cases in Louisiana, USA. Their proposed algorithm combines selecting similar day features using Xgboost, K-Means, and LSTM. In [arora2020prediction], researchers in India compared the Convolutional LSTM, BILSTM, and Stacked LSTM algorithms to the MAPE test to use an optimal algorithm to predict the number of coronary arteries in India Other researchers used simple Recurrent Neural Network (RNN) algorithms, LSTM, two-way LSTM (BiLSTM), Gateway Recurrent Units (GRUs), and Variable Automatic Encoder (VAE) to predict COVID-19 in six countries [zeroual2020deep].
@@
Authors in [kafieh2020isfahan] used time-series data related to data from COVID-19 in Iran. Using the LSTM algorithm, they sought to include the interaction of all classes of coronary arteries, deceased coronary arteries, and the number of recovered individuals in the prediction process using this method. [abbasimehr2021prediction] used the multi-head attention-based method (ATT_BO), CNN-based method (CNN_BO), and LSTM-based method (LSTM_BO) to forecast short-term and long-term COVID-19. For this reason, they perform two types of datasets. [achterberg2020comparing] used a Network Inference based Prediction Algorithm (NIPA) to forecast the corona in the Chinese city of Hobby and the Netherlands. They also compared their method with LSTM and sigmoid curve methods In the same years, [galasso2022random] from the USA presented a relatively non-parametric random forest model to forecast the number of COVID-19 cases in the U.S. [xu2022forecasting] used CNN, LSTM, and the CNN-LSTM algorithm to predict the number of COVID-19 in Brazil, India, and Russia. They also compared the performance of these models with the previously developed deep learning models. [zhou2022interpretable] used the Interpretable Temporal Attention Network (ITANet) algorithm to forecast COVID-19. By using this algorithm, they wanted to infer the importance of government interventions.
@@
Consequently, neural networks and deep learning algorithms proved to be a promising approach that can forecast with high accuracy and are compatible with large data sets and complex systems which have contributed to more than 40 percent of epidemy forecasting [rahimi2021review] and many pieces of research other than those mentioned previously have been devoted to deep learning algorithms [chimmula2020time, ayyoubzadeh2020predicting, fong2020finding, tamang2020forecasting, horby2020effect, sujath2020machine, sholehrasa2024integrating]. Although much research has been developed during the Corona to provide the best predictive results, very few studies have provided a comprehensive analysis of the combination of different neural network algorithms to optimize results and increase prediction accuracy. In other words, to use the important and practical features of different algorithms, they can be combined together to get better results from the models. Also, the use of Metaheuristic algorithms has rarely been investigated in this area. To this end, and to optimize the results, this study, in addition to using a hybrid neural network, for the first time combines two Metaheuristic algorithms of gray wolf and whale on the proposed neural network. Finally, the proposed model is examined on the data of twenty-four countries from six continents in the world.
@@ 2.1 Convolutional Neural Network
During the 1960s, Hubel and his colleagues conducted biological research that demonstrated how visual information is transmitted from the retina to the brain through multiple levels of receptive field excitation. This discovery eventually led to the development of Convolutional Neural Networks (CNNs). A CNN is a type of feed-forward neural network composed of various layers such as input, convolution, pooling, fully connected, and output layers. The input layer processes data through feature transformation and extraction using convolution and pooling layers, which integrate local information. The fully connected layers then map this information to the output signals generated by the output layer after receiving specific properties [lecun2015deep]. The structure of a CNN is depicted in Figure 1, and its calculation formula is presented in Equation 1, where N, W, and F represent the output size, input size, and size of the kernel, respectively. Additionally, P and S denote padding size and step size, respectively.
@@
The main function of the convolution layer is to extract features from the input data. Typically, the convolution layer of a CNN is made up of three main components: the convolution kernel, convolutional layer parameters, and an activation function. This layer is widely regarded as the most essential and standout component within a CNN architecture. The convolution layer in a CNN leverages convolution kernels to extract features from input variables. The size of the kernel employed has usually a smaller value compared to that of the input matrix. Unlike general matrix operations, the computation within convolutional layers utilizes The feature map output generated through convergence operations, wherein each element within the feature map is computed using Equation 2. [wang2017deterministic]. This equation denotes the output value at row $i$ column $j$ of the feature map (referred to as $x_{i,j}^{out}$)), in conjunction with the corresponding value at row $i$ column $j$ of the input matrix (denoted as $x_{i+m,j+n}^{in}$) and the weight (represented as $w_{m,n}$ within row $m$ and column $n$ of the convolution kernel. Additionally, the bias (represented as b) associated with the convolution kernel is taken into consideration, along with the utilization of the chosen activation function (denoted as $f_{cov}(0)$ ).
@@
Generally, the input matrix within the convolutional layer of Convolutional Neural Networks (CNNs) leverages multiple kernels. Each individual kernel operates by extracting distinctive features from the input matrix, subsequently generating a corresponding feature map. Subsequently, the pooling layer assumes the task of reducing the dimensions of the preceding feature map, effectively decreasing both its length and width. This down-sampling operation not only enhances computational efficiency but also contributes to overall performance improvement. The convolutional layer’s output, comprising feature vectors, can be further diminished in dimensionality through the application of the pooling layer. Remarkably, this reduction process simultaneously results in enhanced outcomes. This is attributed to CNNs’ inherent capability to proficiently extract features from grid-like data structures. As a result, a technique involving the expansion of m variables of arbitrary nature into n stations has been employed to obtain a matrix consisting of m rows and n columns. Thus, one can regard the CNN, as a whole, as encompassing a fully connected layer that serves as a classifier. Positioned at the terminal section of the network, this fully connected layer conducts regression classification on the features that have been extracted. Consequently, the CNN can be logically divided into two distinct parts: the initial segment involves feature extraction, encompassing operations such as convolution, activation functions, and pooling. On the other hand, the latter part of the CNN architecture focuses on classification and recognition, primarily facilitated by the fully connected layer [yao2021optimal]. For a visual representation of the CNN model’s structural layout, refer to Figure 1.
@@ 2.2 Long Short-Term Memory
RNNs have gained significant popularity in the fields of time-series analysis and sequence modeling due to their unique ability to incorporate information from past observations through the utilization of directed cycles. This characteristic enables RNNs to effectively process current data by leveraging the knowledge acquired from previous data points. They employ a memory mechanism to remember the state of previous data but they face Vanishing and Exploding Gradient problems when learning long-term dependencies. The LSTM and GRU networks are advanced models of the RNNs, proposed to solve these problems [srivastava2018comparative, ko2020deep, zhang2021deep].
@@
In LSTM, once the recurrent neural network "forgets" a portion of the previous state, the input gate plays a crucial role in incorporating the latest memory from the current input. This process is accomplished through the “input gate”. The input gate in LSTM comprises two components. The first component is a sigmoid layer referred to as the “input threshold layer” which determines the values that need to be updated. The second component is a $tanh$ layer, that generates a new candidate vector $\tilde{C}_{t}$, which is subsequently incorporated into the state. This relationship is illustrated in Equations 4, 5, and 6
@@
In Equations 4, 5, and 6, $W_{n}$ denotes the weight matrix, $b_{n}$ denotes the bias term, $W_{m}$ corresponds to the weight matrix used to update the unit status, $b_{m}$ represents the bias term used to update the unit status [wang2019forecasting], and $C_{t}$ denotes the updated memory unit status. Equation 7 involves the input gate $h_{t}$ and ${C}_{t}$, which undergo a dot product operation to determine whether the state of the time step memory unit should be updated. The forget gate $F_{t}$ performs a scalar product with $C_{t-1}$ to determine whether it is necessary to retain the initial state of the memory unit for the time step.
@@ 2.3 Hyperparameter Tuning
As hyperparameter values play a critical role in the performance of deep learning networks, researchers frequently resort to techniques such as Grid and Random search to identify the most suitable hyperparameter configurations for deep neural networks[alizadeh2020network]. The former is a traditional method that needs a set of manually specified values for each hyper-parameter. Then, it simply explores all the search space to find the optimal values. Even though it can be parallelized easily, the grid search suffers from the curse of dimensionality. On the other hand, a random search can sample the important dimension of search space, so performs more efficiently than a grid search. Employing nature-inspired meta-heuristic algorithms like Genetic Algorithms and Swarm Intelligence is an alternative approach to optimize hyper-parameters of neural networks with promising results. In recent times, metaheuristic algorithms have gained attention for refining hyperparameters[alizadeh2021combination]. In this study, using the combination of the gray wolf optimization algorithm (GWO) and the whale optimization algorithm (WOA) in the form of a random switch is proposed to determine the optimal values of the hyperparameters. The following subsection presents a detailed explanation of this algorithm.

### v3

@@ 1 Introduction
COVID-19 affected the lives of people all over the world and took the lives of thousands around the world. On the other hand, the social and economic consequences of COVID-19 sometimes inflicted irreparable damage on countries, and it was able to confront the world with new challenges and crises. Therefore, many researchers have tried to minimize the damage caused by this disease by providing models to predict the rate of coronary artery disease. Various approaches to forecasting have also been developed. For instance, [22] developed a discrete-time stochastic model to describe the dynamics of pandemic expansion. Prediction Models Traditional time series have been routinely studied to predict COVID-19 cases [29, 66, 73]. However, Stochastic modeling has some drawbacks such as being incapable of handling large data scales, modeling complex nonlinear characteristics, and its dependence on pre-defined features or assumptions[58]. Furthermore, the traditional Autoregressive Integrated Moving Average (ARIMA) model proved to be a suitable approach for describing short-term autocorrelation in time series data, researchers have commonly used it. [12] utilized a time series method such as the ARIMA to predict the number of validated items and compared to other conventional methods such as stochastic modeling, which ARIMA proved to be more accurate in short-time forecasting. Other studies have been developed in India using traditional ARIMA modeling [20]. In 2022, other researchers from Bangladesh used ARIMA to predict verified people, deaths, and recoveries in COVID-19 [41]. [46] used ARIMA, cubist regression (CUBIST), random forest (RF), Ridge Regression (RIDGE), Support Vector Regression (SVR), and stacking-ensemble learning to forecast the total number of patients in Brazil in 1, 3, and 6 days. ARIMA has had superior performance in the case of short-time forecasting. There are also some other studies devoted to developing ARIMA approaches for forecasting the Covid epidemy [48, 30, 52, 37, 34, 15, 6].
@@
While ARIMA may provide a good fit for the initial spread of COVID-19, its effectiveness may diminish as the pandemic progresses and new interventions are implemented [23]. On the other hand, another approach, namely logistic function, may allow for a more flexible and dynamic approach to modeling the pandemic’s impact on a particular population. There are studies on developing a logistic function to predict the spread of COVID-19 [47]. The logistic function method assumes that the spread of COVID-19 can be modeled using a sigmoidal curve, which allows for the prediction of an eventual plateau or decline in the number of cases, which are discussed completely in [9, 32, 43] which proved to be an appropriate approach when the number of cases is decreasing.
@@
In [63] a more advanced algorithm named Logistic Patient Information Based Algorithm (LPIBA) to estimate the number of COVID-19-related deaths in China. Their results show that the overall mortality rate in Hubei and Wuhan is 13 percent and between 0.75 percent and 3 percent in the rest of China. Furthermore, in [16], the LPIBA was developed to address forecasting the infected cases. The LPIBA proved to have superior performance over ARIMA, having features such as Incorporation of patient information, flexibility in terms of data type, and ability to handle binary outcomes. However, it is worth mentioning the LPIBA has limitations when detailed data is not available or it is not feasible to collect some features. [38] used a hybrid ARIMA and Prophet Model to predict daily confirmed and cumulative confirmed cases in India. The Prophet model is open-source developed by Facebook Data and proved to have appropriate performance with time-series types of data having seasonal effects [56, 57]. It proved to be a robust algorithm in the case when there are missing data [39]. In another work, Hosseini et al. proposed a Bayesian predictive method to predict the West Nile virus cases[26], however, these methods also need careful parameter selection and prior distribution assumptions.
@@
Another important and popular approach to epidemic modeling involves treating disease spread as stochastic processes over networks [42]. Historically, these methods were constrained to small populations, typically on the order of a few thousand individuals. However, the recently introduced FastGEMF module has enabled simulations encompassing millions of nodes over multi or single-layer complex networks [25]. While network-based disease spread models provide precise insights into disease propagation, their stochastic nature and the high degree of freedom (DOF) in selecting network parameters or disease models—such as SIS, SIR, and others—often necessitate extensive simulations. This approach is particularly suitable when substantial information is available regarding disease dynamics or population structure, as demonstrated in studies such as [13] that model spill-over using the SIR model over interconnected networks with a case study of the Ebola virus to see how the level of interconnection can lead to an epidemic in human networks. In contrast, for newly emerging diseases like COVID-19, purely data-driven methods are often more appropriate [67].
@@
To deal with more complex data, Neural Networks proved to have superior performance over Prophet Model and ARIMA[8, 49, 59]. In [14], authors used the ANN algorithm to forecast the number of cumulative cases of infection and deaths in Brazil. They also used a substantial mitigation procedure adopted (mandatory use of masks) was experimented as an input to evaluate the improvement in the results. [60] developed a more complex NN, a hybrid algorithm for predicting short-term COVID-19 cases in Louisiana, USA. Their proposed algorithm combines selecting similar day features using Xgboost, K-Means, and LSTM. In [5], researchers in India compared the Convolutional LSTM, BILSTM, and Stacked LSTM algorithms to the MAPE test to use an optimal algorithm to predict the number of coronary arteries in India Other researchers used simple Recurrent Neural Network (RNN) algorithms, LSTM, two-way LSTM (BiLSTM), Gateway Recurrent Units (GRUs), and Variable Automatic Encoder (VAE) to predict COVID-19 in six countries [70].
@@
Authors in [27] used time-series data related to data from COVID-19 in Iran. Using the LSTM algorithm, they sought to include the interaction of all classes of coronary arteries, deceased coronary arteries, and the number of recovered individuals in the prediction process using this method. [1] used the multi-head attention-based method (ATT_BO), CNN-based method (CNN_BO), and LSTM-based method (LSTM_BO) to forecast short-term and long-term COVID-19. For this reason, they perform two types of datasets. [2] used a Network Inference based Prediction Algorithm (NIPA) to forecast the corona in the Chinese city of Hobby and the Netherlands. They also compared their method with LSTM and sigmoid curve methods In the same years, [18] from the USA presented a relatively non-parametric random forest model to forecast the number of COVID-19 cases in the U.S. [68] used CNN, LSTM, and the CNN-LSTM algorithm to predict the number of COVID-19 in Brazil, India, and Russia. They also compared the performance of these models with the previously developed deep learning models. [72] used the Interpretable Temporal Attention Network (ITANet) algorithm to forecast COVID-19. By using this algorithm, they wanted to infer the importance of government interventions.
@@
Consequently, neural networks and deep learning algorithms proved to be a promising approach that can forecast with high accuracy and are compatible with large data sets and complex systems which have contributed to more than 40 percent of epidemy forecasting [44] and many pieces of research other than those mentioned previously have been devoted to deep learning algorithms [10, 7, 17, 55, 24, 54, 51]. Although much research has been developed during the Corona to provide the best predictive results, very few studies have provided a comprehensive analysis of the combination of different neural network algorithms to optimize results and increase prediction accuracy. In other words, to use the important and practical features of different algorithms, they can be combined together to get better results from the models. Also, the use of Metaheuristic algorithms has rarely been investigated in this area. To this end, and to optimize the results, this study, in addition to using a hybrid neural network, for the first time combines two Metaheuristic algorithms of gray wolf and whale on the proposed neural network. Finally, the proposed model is examined on the data of twenty-four countries from six continents in the world.
@@ 2.1 Convolutional Neural Network
During the 1960s, Hubel and his colleagues conducted biological research that demonstrated how visual information is transmitted from the retina to the brain through multiple levels of receptive field excitation. This discovery eventually led to the development of Convolutional Neural Networks (CNNs). A CNN is a type of feed-forward neural network composed of various layers such as input, convolution, pooling, fully connected, and output layers. The input layer processes data through feature transformation and extraction using convolution and pooling layers, which integrate local information. The fully connected layers then map this information to the output signals generated by the output layer after receiving specific properties [31]. The structure of a CNN is depicted in Figure 1, and its calculation formula is presented in Equation 1, where N, W, and F represent the output size, input size, and size of the kernel, respectively. Additionally, P and S denote padding size and step size, respectively.
@@
The main function of the convolution layer is to extract features from the input data. Typically, the convolution layer of a CNN is made up of three main components: the convolution kernel, convolutional layer parameters, and an activation function. This layer is widely regarded as the most essential and standout component within a CNN architecture. The convolution layer in a CNN leverages convolution kernels to extract features from input variables. The size of the kernel employed has usually a smaller value compared to that of the input matrix. Unlike general matrix operations, the computation within convolutional layers utilizes The feature map output generated through convergence operations, wherein each element within the feature map is computed using Equation 2. [62]. This equation denotes the output value at row $i$ column $j$ of the feature map (referred to as $x_{i,j}^{out}$)), in conjunction with the corresponding value at row $i$ column $j$ of the input matrix (denoted as $x_{i+m,j+n}^{in}$) and the weight (represented as $w_{m,n}$ within row $m$ and column $n$ of the convolution kernel. Additionally, the bias (represented as b) associated with the convolution kernel is taken into consideration, along with the utilization of the chosen activation function (denoted as $f_{cov}(0)$ ).
@@
Generally, the input matrix within the convolutional layer of Convolutional Neural Networks (CNNs) leverages multiple kernels. Each individual kernel operates by extracting distinctive features from the input matrix, subsequently generating a corresponding feature map. Subsequently, the pooling layer assumes the task of reducing the dimensions of the preceding feature map, effectively decreasing both its length and width. This down-sampling operation not only enhances computational efficiency but also contributes to overall performance improvement. The convolutional layer’s output, comprising feature vectors, can be further diminished in dimensionality through the application of the pooling layer. Remarkably, this reduction process simultaneously results in enhanced outcomes. This is attributed to CNNs’ inherent capability to proficiently extract features from grid-like data structures. As a result, a technique involving the expansion of m variables of arbitrary nature into n stations has been employed to obtain a matrix consisting of m rows and n columns. Thus, one can regard the CNN, as a whole, as encompassing a fully connected layer that serves as a classifier. Positioned at the terminal section of the network, this fully connected layer conducts regression classification on the features that have been extracted. Consequently, the CNN can be logically divided into two distinct parts: the initial segment involves feature extraction, encompassing operations such as convolution, activation functions, and pooling. On the other hand, the latter part of the CNN architecture focuses on classification and recognition, primarily facilitated by the fully connected layer [69]. For a visual representation of the CNN model’s structural layout, refer to Figure 1.
@@ 2.2 Long Short-Term Memory
RNNs have gained significant popularity in the fields of time-series analysis and sequence modeling due to their unique ability to incorporate information from past observations through the utilization of directed cycles. This characteristic enables RNNs to effectively process current data by leveraging the knowledge acquired from previous data points. They employ a memory mechanism to remember the state of previous data but they face Vanishing and Exploding Gradient problems when learning long-term dependencies. The LSTM and GRU networks are advanced models of the RNNs, proposed to solve these problems [53, 28, 71].
@@
In LSTM, once the recurrent neural network "forgets" a portion of the previous state, the input gate plays a crucial role in incorporating the latest memory from the current input. This process is accomplished through the “input gate”. The input gate in LSTM comprises two components. The first component is a sigmoid layer referred to as the “input threshold layer” which determines the values that need to be updated. The second component is a tanhtanh layer, that generates a new candidate vector $\tilde{C}_{t}$, which is subsequently incorporated into the state. This relationship is illustrated in Equations 4, 5, and 6
@@
In Equations 4, 5, and 6, $W_{n}$ denotes the weight matrix, $b_{n}$ denotes the bias term, $W_{m}$ corresponds to the weight matrix used to update the unit status, $b_{m}$ represents the bias term used to update the unit status [64], and $C_{t}$ denotes the updated memory unit status. Equation 7 involves the input gate $h_{t}$ and ${C}_{t}$, which undergo a dot product operation to determine whether the state of the time step memory unit should be updated. The forget gate $F_{t}$ performs a scalar product with $C_{t-1}$ to determine whether it is necessary to retain the initial state of the memory unit for the time step.
@@ 2.3 Hyperparameter Tuning
As hyperparameter values play a critical role in the performance of deep learning networks, researchers frequently resort to techniques such as Grid and Random search to identify the most suitable hyperparameter configurations for deep neural networks[3]. The former is a traditional method that needs a set of manually specified values for each hyper-parameter. Then, it simply explores all the search space to find the optimal values. Even though it can be parallelized easily, the grid search suffers from the curse of dimensionality. On the other hand, a random search can sample the important dimension of search space, so performs more efficiently than a grid search. Employing nature-inspired meta-heuristic algorithms like Genetic Algorithms and Swarm Intelligence is an alternative approach to optimize hyper-parameters of neural networks with promising results. In recent times, metaheuristic algorithms have gained attention for refining hyperparameters[4]. In this study, using the combination of the gray wolf optimization algorithm (GWO) and the whale optimization algorithm (WOA) in the form of a random switch is proposed to determine the optimal values of the hyperparameters. The following subsection presents a detailed explanation of this algorithm.